OpenStudy (anonymous):

FAN AND MEDAL: Given the equation y - 3 = one half(x + 6) in point-slope form, identify the equation of the same line in standard form. x - 2y = -12 y = one halfx y = one halfx + 6 y = one halfx + 9

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

Please Help?

OpenStudy (ipwnbunnies):

Can you simplify it so only 'y' is remaining on the left side? \[y - 3 = \frac{1}{2}(x + 6)\]

OpenStudy (anonymous):

I don't know. I am unfamiliar with this topic...I don't know how to solve it

OpenStudy (ipwnbunnies):

.-. Do you know simple algebra, distributive property, and such?

OpenStudy (mathmale):

\[ y - 3 = (one~half)(x + 6)\rightarrow y-3=\frac{ 1 }{ 2 }(x+6)\]

OpenStudy (mathmale):

All I did here was to write the problem correctly, mathematically.

OpenStudy (anonymous):

Not really. That what this is. I'm taking Algebra 1 right now...

OpenStudy (anonymous):

Okay. Thank you @mathmale

OpenStudy (anonymous):

do I have to distribute the 1/2 ?

OpenStudy (mathmale):

What you need to do next is to solve \[ y-3=\frac{ 1 }{ 2 }(x+6)\] for y. You could distribute that (1/2) if you want, but I have a suggestion that might involve less work.

OpenStudy (mathmale):

To eliminate that fraction, 1/2, please multiply all terms by 2. Note that (1/2)(x+6) counts as ONE term, not as two terms.

OpenStudy (anonymous):

Okay. so do I add 3 to both sides so that y is by itself?

OpenStudy (mathmale):

You could do that if you prefer to attack the problem in that order.

OpenStudy (mathmale):

Your choice.

OpenStudy (anonymous):

Well, how do you recommend? Because I'm not good at this at all... :(

OpenStudy (mathmale):

You've got to be willing to experiment and to take risks. Why not do this problem both ways? Then you can decide which way YOU prefer.

OpenStudy (anonymous):

Okay. So if I subtract the 3, then would you get: y= 1/2(x+6) - 3 ?

OpenStudy (mathmale):

Kayla, please stop saying derogatory things about yourself such as, "I'm not good at this at all." What you don't know you can most likely learn. Beating upon yourself simply raises obstacles instead of tearing them down. Yes, that equation is what you'd get, EXCEPT that the sign of your 3 must be positive, not negative. Why?

OpenStudy (mathmale):

Actually, you are ADDING 3 to both sides of your original equation, not subtracting 3.

OpenStudy (anonymous):

Oh yes. Because I am adding so: y= 1/2(x+6) +3

OpenStudy (mathmale):

Very nice.

OpenStudy (anonymous):

Then whats the next step?

OpenStudy (mathmale):

I'm going to type your equation in Equation Editor for greater clarity. If you know how to use E. E., great; if not, you could either use parentheses or draw your equation in the Draw utility, below. \[y= 1/2(x+6) +3\rightarrow y= \frac{ 1 }{ 2 }(x+6) +3\]

OpenStudy (mathmale):

Now, Kayla, please multiply each of the 2 terms on the right side of this equation by 2 (to remove the fraction (1/2); also multiply the left side of the equation by 2. What do you get?

OpenStudy (anonymous):

Is that the blue button with the sigma that says equation?

OpenStudy (mathmale):

Yes. Let's go into that a bit later. Please let's solve this equation first. \[2(y=\frac{ 1 }{ 2 }(x+6) ) \rightarrow ??\]

OpenStudy (mathmale):

Multiply each of the two terms within the outer set of parentheses by 2.

OpenStudy (anonymous):

\[(2)y = (2)\frac{ 1 }{ 2 }(x+6) +3(2)\]

OpenStudy (mathmale):

Good. That boils down to 2y on the left side and (what?) on the right side? I see you've found my oversight. You were and are correct to have included that 3(2) (which I'd neglected to write in).

OpenStudy (anonymous):

so:\[2y=2x+12 +6]

OpenStudy (anonymous):

\[2y= 2x +18\]

OpenStudy (anonymous):

right?

OpenStudy (mathmale):

Please look at this again:\[(2)y = (2)\frac{ 1 }{ 2 }(x+6) +3(2).\] Doesn't (2)(1/2) reduce to 1?

OpenStudy (mathmale):

\[(2)y = (2)\frac{ 1 }{ 2 }(x+6) +3(2)\]

OpenStudy (mathmale):

If 2(1/2) reduces to 1, then we get:\[2y=1(x+6)+6.\]

OpenStudy (mathmale):

check this out and convince yourself that this is correct; if you can't, let's discuss it further.

OpenStudy (anonymous):

Wait, wouldn't you multiply the (x+6) by 2 as well?

OpenStudy (mathmale):

You already have done that:\[2(\frac{ 1 }{ 2 }(x+6))=\frac{ 1 }{ 2 }(2)(x+6)=??\]

OpenStudy (anonymous):

Try bringing it to the for Ax+By+C=0...

OpenStudy (anonymous):

Oh okay. So then what do i do? Do i divide everything by 2 so that the y is alone again?

OpenStudy (anonymous):

Explain? @Yajnesh_Joglekar

OpenStudy (anonymous):

Multiply the original eqn by 2...

OpenStudy (anonymous):

Tell me what u get

OpenStudy (mathmale):

Let me ask you this: What is the value of 2(1/2)? of (1/2)(2)?

OpenStudy (anonymous):

2/2 which is 1. And 2/2 which is 1.

OpenStudy (anonymous):

Ans this !... :O

OpenStudy (anonymous):

I continued on the math and I came up with this: x - 2y = -12 is it right?

OpenStudy (anonymous):

U got it.. :P :D

OpenStudy (anonymous):

Awesome! I was kinda confused at first..but I think what you told me about finding a way to better understand it myself helped. @mathmale Thanks to both of you! @mathmale and @Yajnesh_Joglekar

OpenStudy (mathmale):

Let me summarize: \[2(\frac{ 1 }{ 2 }(x+6))=\frac{ 1 }{ 2 }(2)(x+6)\]becomes\[2(\frac{ 1 }{ 2 }(x+6))=\frac{ 1 }{ 2 }(2)(x+6)=x+6. \] So then we end up with 2y=x+6+ 6, or 2y=x+12. Since our job is to solve for y, please divide both sides of this latest equation by 2. What do you get?

OpenStudy (anonymous):

Is the answer I got not right? And @mathmale May you help me a little more?

OpenStudy (mathmale):

Remember, Kayla, your objective here is to SOLVE FOR y. Your present result is x-2y=-12, which has not yet been solved for y.

OpenStudy (anonymous):

Not necessarily. It's asking for me to identify the equation of the same line in standard form.

OpenStudy (mathmale):

Much of our discussion here has consisted of reviews. the actual problem solution is pretty straightforward, once you have the experience: Actually, you're doing better than I here. I've overlooked that "standard form" until now. Your equation, x-2y=-12 IS already in standard form. sorry about that. I'm going to double-check your result quickly, but at first glance I believe it's correct. What do y ou want to do next? if you have further questions, please post them, one at a time in the "Ask a question" box.

OpenStudy (anonymous):

Okay. Thank you

OpenStudy (mathmale):

Your result is perfect. congrats!

OpenStudy (anonymous):

Both you and @Yajnesh_Joglekar have helped me tremendously. I think i'm beginning to understand this better. And thanks @mathmale ! I wouldn't have been able to find the answer without you! (:

OpenStudy (mathmale):

To demo how brief the actual problem solution can be:\[y-3=\frac{ 1 }{ 2 }(x+6)\]Multiply both sides by 2. Result:\[2(y-3)=x+6\]

OpenStudy (anonymous):

Haha Wow. I could have been done forever ago lol And I tagged you in a problem

OpenStudy (mathmale):

Distribute that 2 over (y-3):\[2y-6=x+6\] Subtract 2y from both sides of this equation, and then subtract 6 from both sides:\[2y-6\] (I'm not finished yet, but if you understand this process and already have the correct answer, let's move on!

OpenStudy (anonymous):

I'm good hun (: